# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v3

@@ Abstract
Large Language Model (LLM) alignment objectives are often presented as a collection of distinct algorithms—PPO, DPO, IPO, and their variants—each motivated by different derivations. In this work, we argue that this diversity obscures a simpler underlying structure. At a fundamental level, alignment objectives involve two independent design choices: (i) how training signals are sampled and weighted, and (ii) how deviations from a reference policy are geometrically penalized. Existing methods typically entangle these choices through a single divergence, most commonly the Kullback–Leibler (KL) divergence.
We show that this entanglement is not merely a modeling convenience but a source of systematic instability. When the same divergence simultaneously determines sample weighting and optimization curvature, adjusting one aspect (e.g., exploration strength) inevitably alters the other (e.g., gradient geometry). This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
We propose a principled structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters as a linear driving force derived from an $\alpha$-divergence projection, while optimization geometry is determined independently by a Bregman divergence (mirror map). This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space. The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
@@ 1 Introduction
While empirically effective, KL-regularized objectives exhibit distinct limitations in reasoning-intensive domains where high confidence is required. As the policy improves and assigns high probability to correct reasoning chains, the KL penalty—which induces an exponential geometry in the log-probability space—often dominates the learning signal. This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident. Consequently, training often plateaus prematurely, behavior commonly attributed to "over-regularization" but which we argue is intrinsic to the chosen geometry.
@@
To resolve this, we propose a unifying framework based on Orthogonal Mirror Descent. By viewing alignment as a mirror descent process, we can explicitly decouple these two axes: the sampling geometry acts as a linear driving force defined by an advantage-weighted distribution, while the optimization geometry is independently governed by a Bregman divergence (mirror map). This perspective reveals that the saturation observed in DPO and PPO is not a necessary feature of alignment, but a specific consequence of using the Negative Entropy mirror map (which induces KL divergence).
Guided by this framework, we introduce Orthogonalized Policy Optimization (OPO), a new alignment objective derived by selecting the Euclidean mirror map in the space of likelihood ratios. Unlike the exponential geometry of KL, the Euclidean geometry induces a Pearson $\chi^{2}$ trust region, leading to a quadratic regularizer that maintains a linear, non-saturating driving force even in high-confidence regimes. This formulation cleanly separates the sampling intensity (controlled by an $\alpha$-divergence parameter) from the regularization strength (controlled by a stiffness parameter), allowing for robust, continuous improvement.
@@
  • We propose the Orthogonal Mirror Descent framework, which decouples these design axes via two independent geometries: an $\alpha$-divergence projection governs sampling, while a Bregman divergence (mirror map) governs optimization.
  • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
  • We empirically demonstrate that OPO outperforms strong baselines like DPO and GRPO on mathematical reasoning tasks, validating the benefits of non-saturating gradients.
@@ 2 Related Work
3 Preliminaries: Two Coordinate Systems
Before presenting the OPO framework, we establish the two coordinate systems that underlie all ratio-based alignment objectives.
3.1 Ratio Space vs. Log-Ratio Space
Let the reference policy be $\pi_{\mathrm{ref}}$. We define two coordinate systems for measuring policy deviation:
Definition 3.1 (Ratio Coordinates).
The density ratio and centered ratio are:
$t_{\theta}(y):=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)},\qquad v_{\theta}(y):=t_{\theta}(y)-1=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
Definition 3.2 (Log-Ratio Coordinates).
The log-ratio (natural parameterization of LLMs) is:
$\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ (2)
so that $t_{\theta}(y)=\exp(\Delta_{\theta}(y))$ and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$.
The key observation is that $\chi^{2}$ divergence is naturally quadratic in ratio coordinates $v$, while KL divergence involves the log-ratio $\Delta$. This distinction is crucial for understanding the gradient dynamics of different alignment objectives.
4 Theoretical Framework: Orthogonal Mirror Descent
We present a unified framework that fundamentally decouples the preference sampling geometry (what to learn) from the optimization geometry (how to learn). We formulate policy alignment as a functional Orthogonal Mirror Descent problem.
4.1 The General Framework
We view alignment as an iterative projection process. At each step $k$, we seek a policy $\pi$ that maximizes the expected return under a specific sampling distribution $\rho_{\alpha}$, subject to a trust-region constraint defined by a divergence $D_{\Psi}$:
$\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{y\sim\rho_{\alpha}}\left[A(y)\log\pi(y)\right]-\frac{1}{\eta}D_{\Psi}(\pi\|\pi_{k})$ (3)
This formulation cleanly orthogonizes two design axes:
  1. 1.
    Sampling Geometry ($\rho_{\alpha}$): Determines the effective weighting of samples (the linear driving force).
  2. 2.
    Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
4.2 Axis 1: Sampling Geometry via $\alpha$-Divergence Projection
The sampling weight $\omega_{\alpha}$ determines the effective weighting of training examples. We derive it from a principled $\alpha$-divergence interpolation, providing the sampling axis with the same geometric rigor as the optimization axis.
Definition 4.1 (Oracle Target Distribution).
Given a non-negative quality signal $Q(y)\geq 0$ and the behavior policy $\tilde{p}(y):=\pi_{\mathrm{old}}(y)$, define the oracle target distribution:
$P^{*}(y)\;\propto\;Q(y)$ (4)
In the standard RL setting with advantage function $A(y)$, the natural choice is $Q(y)=\exp(A(y))$, corresponding to the Boltzmann-optimal density. The importance sampling (IS) ratio is $w_{\mathrm{IS}}(y)=P^{*}(y)/\tilde{p}(y)\propto Q(y)/\tilde{p}(y)$.
Direct use of $w_{\mathrm{IS}}$ suffers from variance explosion in large action spaces. The parameter $\alpha\in[0,1]$ geometrically tempers this correction by interpolating along the $\alpha$-geodesic [11] connecting $\tilde{p}$ to $P^{*}$:
Proposition 4.2 ($\alpha$-Escort Weighting).
The $\alpha$-geodesic between $\tilde{p}$ and $P^{*}$ yields the escort distribution $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$. The corresponding importance weight is:
$\boxed{\omega_{\alpha}(y)\;=\;\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\;\propto\;\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}=\left(\frac{\exp(A(y))}{\tilde{p}(y)}\right)^{1-\alpha}}$ (5)
This parameterization has clear geometric semantics:
  • $\alpha\to 1$: $\omega_{\alpha}\to 1$ (uniform weighting; mode-covering, conservative).
  • $\alpha\to 0$: $\omega_{\alpha}\to P^{*}/\tilde{p}$ (full IS correction; peak-seeking, aggressive).
  • $0<\alpha<1$: Smooth interpolation governed by $\alpha$-geometry.
Remark 4.3 (On-Policy Implementation with Frozen Denominator).
In the on-policy setting where $\tilde{p}=\pi_{\mathrm{old}}$, the theoretical escort weight $\omega_{\alpha}=(\exp(A)/\tilde{p})^{1-\alpha}$ can be decomposed as:
$\omega_{\alpha}(y)\;=\;\exp\!\big((1-\alpha)\,A(y)\big)\;\cdot\;\operatorname{sg}\!\left(\frac{1}{\pi_{\mathrm{old}}(y)^{\,1-\alpha}}\right)$ (6)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. This formulation is not an approximation—it computes the exact escort weight from Eq. (5). The stop-gradient on the $1/\pi_{\mathrm{old}}^{1-\alpha}$ factor serves two critical purposes: (i) it prevents the $1/p$ denominator from injecting unbounded gradients, preserving the constant Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$; and (ii) it is semantically natural, since $\pi_{\mathrm{old}}$ is already frozen as the reference policy in each iteration.
Remark 4.4 (Two-Geometry Structure).
With this derivation, the entire OPO framework admits a clean geometric decomposition:
Alignment == $\alpha$-geometry (sampling) $\times\;$ mirror geometry (optimization)
The sampling axis is an $\alpha$-divergence interpolation in the statistical manifold; the optimization axis is a Bregman divergence projection in ratio space. These two geometries are structurally orthogonal: $\alpha$ shapes the driving force without altering the Hessian, and $\mu$ shapes the curvature without altering the target.
4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
The core innovation of OPO is the choice of the mirror map. Instead of the standard Negative Entropy map (which leads to KL divergence and exponential updates), we choose the Euclidean Mirror Map in the space of likelihood ratios $v=\pi/\pi_{\mathrm{ref}}-1$:
$\Psi(v)=\frac{1}{2}\|v\|_{L^{2}(\pi_{\mathrm{ref}})}^{2}=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}[v(y)^{2}]$ (7)
The induced Bregman divergence is the Pearson $\chi^{2}$ divergence:
$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (8)
This choice transforms the alignment problem into a quadratic optimization in ratio space, with curvature $\mu I$ independent of the data.
4.4 The OPO Objective
Combining the $\alpha$-escort sampling weight (Axis 1) with the Euclidean mirror map (Axis 2) yields the OPO objective:
@@ 4.4 The OPO Objective
$\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (9)
where $\mu=1/\eta$ is the stiffness parameter, with closed-form solution $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
4.5 Gradient Dynamics: Why OPO Does Not Saturate
Differentiating Eq. (9) w.r.t. $v_{\theta}(y)$:
$\nabla_{v}\mathcal{L}_{\text{OPO}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (10)
with unique equilibrium $v^{*}(y)=\omega_{\alpha}(y)/\mu$. This linear relationship has three desirable properties:
  1. 1.
    Stable Equilibrium: Strict convexity in $v_{\theta}$ yields a unique stationary point.
  2. 2.
    No Saturation: Unlike sigmoid-based losses, the gradient does not vanish as $v$ grows—it increases proportionally to the distance from equilibrium.
  3. 3.
    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the curvature. The two do not interact.
4.6 Implementation: Log-Ratio Approximation
The ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$ is not directly available from LLM logits. However, for the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ (which is directly computable), the first-order Taylor expansion $v=e^{\Delta}-1\approx\Delta$ holds within the trust region. This yields the implementable form:
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (11)
This retains the linear dynamics of Euclidean geometry while operating entirely in log-space. The approximation is justified by on-policy anchoring: setting $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ ensures $\Delta\approx 0$ at the start of each iteration.
4.7 Constructive Derivation: Regression in Ratio Space
We provide an alternative, constructive derivation that arrives at the same OPO loss from pure first principles—no mirror maps required.
Step 1: Define the target offset.
Given the escort weight $\omega_{\alpha}$ and regularization strength $\mu$, the optimal ratio offset is $v^{*}(y):=\omega_{\alpha}(y)/\mu$. This is the ideal deviation from the reference: the quality signal drives the direction; $\mu$ bounds the magnitude.
Step 2: Minimize the distance.
The natural metric in ratio space is mean squared error:
$\mathcal{J}(\theta)=\frac{\mu}{2}\big(v_{\theta}(y)-v^{*}(y)\big)^{2}$ (12)
Step 3: Complete the square.
Expanding and discarding the $\theta$-independent constant $\omega_{\alpha}^{2}/(2\mu)$:
$\mathcal{J}(\theta)=\underbrace{\frac{\mu}{2}\,v_{\theta}^{2}}_{\text{quadratic regularizer}}-\underbrace{\omega_{\alpha}\cdot v_{\theta}}_{\text{linear drive}}+\text{const}\;\;\Longrightarrow\;\;\mathcal{L}(\theta)=-\omega_{\alpha}\,v_{\theta}+\frac{\mu}{2}\,v_{\theta}^{2}$ (13)
which is exactly the OPO objective (Eq. 9). The completing-the-square identity provides a constructive proof: MSE regression in ratio space is gradient-equivalent to the linear-plus-quadratic OPO loss.
Remark 4.5 (Equivalence of Perspectives).
[…]
Remark 4.6 (Analogy with Quantum Amplitudes).
A physical analogy: in quantum mechanics, superpositions are linear in amplitude ($\psi$), while measurements are quadratic in amplitude. Similarly, OPO operates in the ratio space $v$ (“amplitude”), where preference signals superpose linearly, and regularization is quadratic. KL-based methods operate in log-probability (“energy”) space, where strong signals lead to exponential collapse. The amplitude picture explains why OPO sustains non-saturating gradients in high-confidence regimes.
5 Theoretical Analysis
The preceding section derived OPO from two complementary perspectives (mirror descent and ratio regression). We now prove that OPO’s key structural properties—orthogonal control and global contraction—hold rigorously.
5.1 Formal Decoupling of Geometry Axes
All analysis is conducted in the function space of ratio outputs $v\in L^{2}(\pi_{\mathrm{ref}})$, which captures the intrinsic objective geometry. We first bridge to parameter space.
Lemma 5.1 (Parameter Space Gradient Projection).
Let $\mathcal{L}(v_{\theta})$ be a functional of the policy output $v_{\theta}(y)$. The gradient in parameter space $\theta$ is the projection of the function-space gradient $\nabla_{v}\mathcal{L}$ onto the tangent space of the policy manifold:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\nabla_{v}\mathcal{L}(v_{\theta}(y))\cdot\nabla_{\theta}v_{\theta}(y)\right]$ (14)
Thus, the geometric properties of $\nabla_{v}\mathcal{L}$ directly determine the driving force in parameter space, modulated by the policy’s local sensitivity $\nabla_{\theta}v_{\theta}$.
With this link established, we characterize the geometry of $\mathcal{L}(v)$.
Theorem 5.2 (Decoupled Sampling and Optimization Geometry).
Consider the OPO objective $\mathcal{L}(v)=-\sum_{y}\omega_{\alpha}(y)v(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]$. In the space of ratio functions $v\in L^{2}(\pi_{\mathrm{ref}})$:
  1. 1.
    The first-order driving force depends on the sampling parameter $\alpha$:
    $\nabla_{v}\mathcal{L}=-\omega_{\alpha}(y)+\mu v(y)$ (15)
  2. 2.
    The second-order curvature is independent of $\alpha$:
    $\nabla^{2}_{v}\mathcal{L}=\mu I$ (16)

### v4

@@ Abstract
We present Orthogonalized Policy Optimization (OPO), a unified theoretical account of LLM alignment grounded in a work–dissipation principle. The policy update is characterized as a constrained proximal response that maximizes external work induced by an $\alpha$-escort sampling field, while paying an intrinsic dissipation cost given by a quadratic fluctuation energy in $\chi^{2}$ (ratio) geometry. This single variational principle simultaneously admits three equivalent interpretations: (i) a mirror-descent step with a Euclidean mirror map in ratio space, (ii) a Hilbert-space projection via the orthogonal projection theorem in $L^{2}(\pi_{k})$, and (iii) a linear-response law from near-equilibrium statistical mechanics. Their convergence to the same closed-form update confirms that OPO is the unique quadratic proximal response within ratio geometry. The framework cleanly decouples sampling geometry ($\alpha$) from optimization geometry ($\mu$), yields constant Hessian and non-saturating linear gradients, and reveals that advantage $z$-score normalization is not a heuristic but a conservation-law projection. Experiments on mathematical reasoning tasks demonstrate that OPO outperforms GRPO, GSPO, and DAPO while maintaining healthy gradient dynamics throughout training.
@@ 1 Introduction
While empirically effective, KL-regularized objectives exhibit distinct limitations in reasoning-intensive domains where high confidence is required. As the policy improves and assigns high probability to correct reasoning chains, the KL penalty—which induces an exponential geometry in the log-probability space—often dominates the learning signal. This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident. Consequently, training often plateaus prematurely, behavior commonly attributed to “over-regularization” but which we argue is intrinsic to the chosen geometry.
@@
To resolve this, we propose a unifying framework based on a work–dissipation principle: the policy update maximizes external work induced by an $\alpha$-escort sampling field, while paying an intrinsic dissipation cost in $\chi^{2}$ ratio geometry. This single variational principle cleanly decouples the two axes and admits three equivalent mathematical interpretations—mirror descent, Hilbert-space projection, and linear response—each illuminating a different structural facet.
@@
  • We propose a unified work–dissipation principle that decouples these design axes: $\alpha$-escort shapes the driving force, while a $\chi^{2}$ quadratic penalty shapes the optimization curvature.
  • We show that the resulting OPO objective admits three equivalent interpretations—Bregman mirror descent, Hilbert-space orthogonal projection, and near-equilibrium linear response—confirming it is the unique quadratic proximal response in ratio geometry. The Hilbert derivation further reveals that advantage $z$-score normalization is a conservation-law projection.
  • We empirically demonstrate that OPO outperforms strong baselines (GRPO, GSPO, DAPO) on mathematical reasoning tasks. A gradient response efficiency analysis—normalizing gradient magnitude by the advantage signal amplitude—reveals that OPO maintains ${\sim}5{\times}$ higher gradient-per-signal than KL-based baselines, providing direct quantitative evidence for the linear response prediction.
@@ 2 Related Work
3 Preliminaries: Ratio Coordinates and Conservation
Let $\pi_{k}$ denote the anchor (reference) policy at iteration $k$, with $\pi_{k}(y)>0$ on its support.
Definition 3.1 (Ratio and Log-Ratio Coordinates).
Define
$t(y):=\frac{\pi(y)}{\pi_{k}(y)},\qquad v(y):=t(y)-1=\frac{\pi(y)}{\pi_{k}(y)}-1,\qquad\Delta(y):=\log\pi(y)-\log\pi_{k}(y).$ (1)
Then $t(y)=e^{\Delta(y)}$ and $v(y)=e^{\Delta(y)}-1$.
Conservation law.
Probability normalization $\sum_{y}\pi(y)=1$ is equivalent to
$\mathbb{E}_{\pi_{k}}[t]=1\quad\Longleftrightarrow\quad\mathbb{E}_{\pi_{k}}[v]=0.$ (2)
Hence admissible fluctuations lie in the closed subspace
$\mathcal{H}_{0}:=\{v\in L^{2}(\pi_{k}):\mathbb{E}_{\pi_{k}}[v]=0\}.$ (3)
4 Unified Work–Dissipation Principle
This section provides the unifying core: OPO is a constrained proximal response that maximizes external work induced by sampling geometry, while paying an intrinsic dissipation (fluctuation) energy that stabilizes the update.
4.1 Sampling Geometry as an $\alpha$-Escort Field
Let $P^{*}$ be an oracle target distribution encoding preference/quality. A canonical choice is $P^{*}(y)\propto\exp(A(y))$ where $A$ is an advantage-like signal.
Definition 4.1 ($\alpha$-Escort Distribution and Escort Weight).
For $\alpha\in[0,1]$, define the escort distribution
$\rho_{\alpha}(y)\propto\pi_{k}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}.$ (4)
Define the escort weight (Radon–Nikodym derivative)
$\omega_{\alpha}(y):=\frac{\rho_{\alpha}(y)}{\pi_{k}(y)}.$ (5)
Since $\rho_{\alpha}$ is normalized,
$\mathbb{E}_{\pi_{k}}[\omega_{\alpha}]=1.$ (6)
Proposition 4.2 ($\alpha$-Geometric Interpolation).
The family $\{\rho_{\alpha}\}_{\alpha\in[0,1]}$ forms an $e$-geodesic connecting $P^{*}$ (at $\alpha=0$) to $\pi_{k}$ (at $\alpha=1$). The escort weight simplifies to $\omega_{\alpha}(y)\propto\bigl(\exp(A(y))/\pi_{k}(y)\bigr)^{1-\alpha}$.
Geometric semantics.
$\alpha$ controls how aggressively sampling emphasizes high-quality regions: $\alpha\to 1$ gives $\omega_{\alpha}\to 1$ (conservative); $\alpha\to 0$ gives $\omega_{\alpha}\to P^{*}/\pi_{k}$ (aggressive).
On-policy implementation (stop-gradient).
When $\pi_{k}=\pi_{\mathrm{old}}$ is frozen during an iteration,
$\omega_{\alpha}(y)\;\propto\;\exp\!\big((1-\alpha)A(y)\big)\cdot\operatorname{sg}\!\Big(\pi_{k}(y)^{-(1-\alpha)}\Big).$ (7)
Batch normalization of escort weights.
The escort density $\rho_{\alpha}(y)\propto\pi_{k}(y)^{\alpha}P^{*}(y)^{1-\alpha}$ involves an unknown normalizing constant $Z_{\alpha}$. In practice we do not compute $Z_{\alpha}$ explicitly; instead, we apply batch-wise zero-centering and $z$-scoring of $\omega_{\alpha}$ over each mini-batch. This is equivalent to estimating the normalizer locally within the sample batch, and ensures that the estimated weights satisfy the mean constraint $\mathbb{E}[\omega_{\alpha}]=1$ at the sample level. The subsequent mean subtraction in the OPO update (Remark 6.2) absorbs any residual bias.
4.2 Intrinsic Dissipation Energy as Quadratic Fluctuation Cost
We stabilize the update by penalizing the fluctuation amplitude $v=\pi/\pi_{k}-1$ in $L^{2}(\pi_{k})$:
$\mathcal{E}_{\mathrm{diss}}(v):=\frac{\mu}{2}\|v\|_{L^{2}(\pi_{k})}^{2}=\frac{\mu}{2}\mathbb{E}_{\pi_{k}}[v(y)^{2}].$ (8)
This is exactly the Pearson $\chi^{2}$ geometry in ratio coordinates: $\mathbb{E}_{\pi_{k}}[v^{2}]=\mathbb{E}_{\pi_{k}}\!\left[(\pi/\pi_{k}-1)^{2}\right]$.
Physical motivation.
Near the anchor, a quadratic fluctuation cost matches the universal Gaussian-fluctuation principle from near-equilibrium thermodynamics: small departures from equilibrium incur a quadratic free-energy cost [19].
4.3 OPO as a Constrained Work–Dissipation Balance
Define the work–dissipation functional
$v^{*}=\operatorname*{arg\,max}_{v\in L^{2}(\pi_{k})}\left\{\underbrace{\langle\omega_{\alpha},v\rangle}_{\text{external work}}\;-\;\underbrace{\frac{\mu}{2}\|v\|^{2}}_{\text{intrinsic dissipation}}\right\}\quad\text{s.t.}\quad\mathbb{E}_{\pi_{k}}[v]=0.$ (9)
Equivalently, in loss-minimization form,
@@ 4.3 OPO as a Constrained Work–Dissipation Balance
$\boxed{\mathcal{L}_{\mathrm{OPO}}(v)=-\mathbb{E}_{\pi_{k}}[\omega_{\alpha}(y)\,v(y)]+\frac{\mu}{2}\mathbb{E}_{\pi_{k}}[v(y)^{2}],\qquad\mathbb{E}_{\pi_{k}}[v]=0.}$ (10)
Orthogonality of the two geometries.
$\alpha$ enters only through $\omega_{\alpha}$ (first-order drive); $\mu$ enters only through the quadratic curvature (second-order geometry). Sampling and optimization are structurally orthogonal.
4.4 Closed-Form Solution (Linear Response Law)
Theorem 4.3 (Closed-Form Update).
The unique optimizer of Eq. (9) is
$\boxed{v^{*}(y)=\frac{1}{\mu}\Big(\omega_{\alpha}(y)-1\Big).}$ (11)